OpenStudy (kittiwitti1):

< closed I guess? > Two Questions - (1) S̶h̶o̶w̶ ̶t̶h̶a̶t̶ ̶t̶h̶e̶ ̶e̶q̶u̶a̶t̶i̶o̶n̶ ̶s̶i̶n̶x̶-̶6̶x̶=̶0̶ ̶h̶a̶s̶ ̶e̶x̶a̶c̶t̶l̶y̶ ̶o̶n̶e̶ ̶r̶o̶o̶t̶ ̶u̶s̶i̶n̶g̶ ̶t̶h̶e̶ ̶M̶e̶a̶n̶ ̶V̶a̶l̶u̶e̶ ̶T̶h̶e̶o̶r̶e̶m̶.̶.̶.̶ I̶ ̶d̶o̶n̶'̶t̶ ̶k̶n̶o̶w̶ ̶h̶o̶w̶ ̶t̶o̶ ̶a̶p̶p̶l̶y̶ ̶t̶h̶e̶ ̶M̶V̶T̶?̶ (2) Find the following: a. intervals of increase/decrease b. local min/max values c. intervals of concavity and inflection points function: f(x)=ln(x^4+1)

OpenStudy (kittiwitti1):

LaTeX: (1) \(\sin{(x)}-6x=0\) (2) \(f(x)=\ln{(x^{4}+1})\)

OpenStudy (tkhunny):

The idea is to: 1) Note that the function is continuous and differentiable in the Domain where you care. 2) Find a value of x such that sin(x) - 6x > 0 3) Find a value of x such that sin(x) - 6x < 0 4) The Mean Value Theorem (MVT) tells us that there must be a value where sin(x) - 6x = 0. However, there is no guarantee there is ONLY one such value. I'd check on the specific wording of the question.

OpenStudy (kittiwitti1):

Er... I don't understand what step/part 1 means, exactly. How would I be able to tell that it's differentiable on an interval...? Or is that what steps 2 and 3 are explaining?

OpenStudy (tkhunny):

You tell me. Is it differentiable? Is ti continuous?

OpenStudy (kittiwitti1):

I don't understand the question...

OpenStudy (tkhunny):

f(x) = sin(x) - 6x Is there a derivative on some interval? Is there anywhere where it jumps around or is it well-behaved?

OpenStudy (kittiwitti1):

I understand continuity but not the derivative part* Sorry.

OpenStudy (tkhunny):

How are you supposed to solve Problem #2? Are you studying the Derivative?

OpenStudy (kittiwitti1):

That problem is from a section titled "How Derivatives Affect the Shape of a Graph"

OpenStudy (kittiwitti1):

I get part (a) but the rest I am not sure if I need to apply certain concepts...

OpenStudy (tkhunny):

Okay, then you should have some idea what a derivative is. Does f(x) = sin(x) - 6x have a derivative?

OpenStudy (kittiwitti1):

I think that it's \(f'(x)=\cos{(x)}-6\)?

OpenStudy (tkhunny):

There you go. Now, on to step #2.

OpenStudy (kittiwitti1):

The first problem is from "Mean Value Theorem" The second problem is from "How Derivatives Affect the Shape of a Graph"

OpenStudy (kittiwitti1):

Okay

OpenStudy (kittiwitti1):

Are steps 2 and 3 variable? (any number that makes said equation true)*

OpenStudy (tkhunny):

Any value will do.

OpenStudy (kittiwitti1):

Alright, thank you. I will work on that... might take a while

OpenStudy (tkhunny):

Nah. Try x = +1 and x = -1.

OpenStudy (kittiwitti1):

Oh... alright lol Just asking... is the MVT related to the IVT in any way?

OpenStudy (tkhunny):

The are similar, but not to be confused.

OpenStudy (kittiwitti1):

Okay, so I got \(f(1)\approx-5.157529...\) and \(f(-1)\approx5.15829...\)

OpenStudy (danjs):

ok, so you know the function f(x)=sin(x) - 6x , is continuous everywhere, and goes from a negative value to a positive value, so it must cross f(x)=0 somewhere At Least Once between -1 and 1. From the intermediate value theorem, you can say there is At least one root inside the interval [-1,1]. That doesn't say that there cant be more than one though...

OpenStudy (danjs):

You can show if there is ONLY one root by assuming there are more and testing to see if it is true. Assume there are 2 roots, and apply the mean value theorem...

OpenStudy (danjs):

If you have 2 roots where f(x)=0, say x=a and x=b, then the mean value theorem will say that there must be some value between a and b where the derivative is zero... \[\large f'(c)=\frac{ f(b)-f(a) }{ b-a }\] If f(a) and f(b) are both roots then \[f'(c)=\frac{ 0-0 }{ b-a }=0\] The function is continuous, so it must turn around somewhere in order to get 2 roots...

OpenStudy (danjs):

Notice the derivative.. f '(x) = cos(x) - 6 For two roots to exist for f(x)=sin(x)-6x , the derivative must be zero at least once.. cos(x) - 6 = 0 That never happens, cos(x)=6, not true. This says that there aren't 2 or more roots for the function.

OpenStudy (danjs):

Overall... From Intermediate value theorem, shown AT Least one Root exists. From Mean Value Theorem, shown impossible for 2 or more roots to exist. From those two, you can say exactly one root Exists.

OpenStudy (kittiwitti1):

@LunaCat please refrain from posting links to your own questions on other people's posts! Thank you ☻

OpenStudy (mathmale):

Could be wrong, but I have the strong feeling that you're discussing the Intermediate Value Theorem here, NOT the Mean Value Theorem. Please double-check on this. "Is f(x) differentiable?" is not a step, but rather a CONDITION. Your function must satisfy at least two conditions before you can even consider applying the IVT or the MVT.

OpenStudy (kittiwitti1):

@mathmale to whom are you referring? Sorry, I am confused.

OpenStudy (issimplcalcus):

(2) Find the following: a. intervals of increase/decrease b. local min/max values c. intervals of concavity and inflection points a. where f'(x) > 0 b. where f'(x) = 0 c. where f''(x) = 0 (inflection) and where f''(x) < or > 0

OpenStudy (kittiwitti1):

Okay, thank you @issimplcalcus ! :)

OpenStudy (mathmale):

I am referring to you, kittiwitti1.

OpenStudy (kittiwitti1):

The question is in the section "Mean Value Theorem" @mathmale

OpenStudy (mathmale):

The MVT and the IVT are usually discussed in the same section. The MVT is, in my opinion, the more important of the two. It doesn't surprise me if you find a problem in this MVT section that asks you to show that there's only one root for the given fn. on the given interval. But the IVT is the relevant theorem here, NOT the MVT.

OpenStudy (loser66):

I would like to see the problem on the new post. It is hard to read a long comment to follow what is going on.

OpenStudy (kittiwitti1):

@Loser66 "Show that the equation \(\sin{(x)}-6x=0\) has exactly one root (using MVT)"

OpenStudy (kittiwitti1):

I'm not quite sure how exactly he wants it, excluding the fact that he needs it done using MVT... :(

OpenStudy (issimplcalcus):

Im dumb whoops

OpenStudy (issimplcalcus):

1. MVT to show a zero exists 2. show (dy/dx) is < 0 for all x

OpenStudy (kittiwitti1):

eh? O-o

OpenStudy (issimplcalcus):

1. MVT to show a zero exists 2. Show that f(x) is only decreasing by showing f'(x) is < 0 for all x

OpenStudy (kittiwitti1):

OHH, I see. o:

OpenStudy (loser66):

You need one value of x which makes f(x) >0 And f(x) monotonic decreases as what issimplcalcus says

OpenStudy (kittiwitti1):

Alright, thank you both! I will work with that in mind. :-)